OpenStudy (anonymous):

convert: 30 mi/hr into ft/ min. (hint: 1 hr= 60 min and 1 mi =5280ft)

OpenStudy (anonymous):

I just answered this exact same question lol ;D I mile = 5,280 ft so multiply that by 30 because there are 30 miles 5,280*30 = 158,400 and there are 60 minutes in an hour so put that over 60 158,400/60 = ? Now just divide that and you have got ft/per min.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

d^2 - 10d +25=6